Please help me figure out:
a) f(x+h) where f(x)=4x+3
I know final answer is 4x+4h-3
But I don't know the steps!

Answer by Boreal(15235)

substitute (x+h) for x
f(x-h)=4(x+h)+3
distribute
=4x+4h+3
Note: there should not be a -3 but a +3. The constant doesn't change.
The next step is to subtract f(x) from it, which is 4x+4h+3-(4x+3)=4h.
Then ultimately you will divide by h to get 4, and that is the derivative of the function.